Nippon Steel’s expenses for heating and cooling a large manufacturing facility are expected to increase according to an arithmetic gradient beginning in year 2. If the cost is $550,000 this year (year 0) and will be $550,000 again in year 1, but then it is estimated to increase by $58,000 each year through year 12, what is the equivalent annual worth in years 1 to 12 of these energy costs at an interest rate of 14% per year? The equivalent annual worth is determined to be $
Answer #1

We need to calculate Equivalent Annual worth in years 1 to 12 which is

= r(NPV)/{1-(1+r)^-n}

Year Cost Discount rate @ 14% NPV
1 550000 0.877193 482456.1404
2 608000 0.769468 467836.2573
3 666000 0.674972 449531.0298
4 724000 0.59208 428666.1208
5 782000 0.519369 406146.2955
6 840000 0.455587 382692.7001
7 898000 0.399637 358874.3156
8 956000 0.350559 335134.4564
9 1014000 0.307508 311813.0541
10 1072000 0.269744 289165.3638
11 1130000 0.236617 267377.6358
12 1188000 0.207559 246580.2137
Total 4426273.583

Equivalent Annual worth = 0.14(4426273.583)/[1-(1+0.14)^-12]

= 619678.3017/(1-0.207559)

= 619678.3017/0.792441

= $ 781986.7747